8.a certain car is capable of accelerating at a rate of 0.57 m/s2. How long does it take for this car to go from a speed of 30 mi/h to a speed of 40 mi/h?

 s


1
Expert's answer
2022-06-06T08:54:51-0400

"30\:\rm mi/h=13.4\:\rm m/s"

"40\:\rm mi/h=17.9\:\rm m/s"


"a=\frac{\Delta v}{\Delta t}"

"\Delta t=\frac{\Delta v}{a}=\frac{17.9-13.4}{0.57}=7.9\:\rm s"